OpenStudy (anonymous):

Explain how to graph x + 3y = 9 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

if you write your equation in the form y=mx+c you would have a - slope. to find x-intercept, y=0 x+3(0)=9 x=9 to find y-intercept, x=0 0+3y=9 y=3